can you sole this please: sqrt X+4 - sqrt X-1 = 2/sqrt x-1
Answer by mananth(16946)

sqrt X+4 - sqrt X-1 = 2/sqrt x-1
..
sqrt(x+4)*sqrt(x-1)-sqrt(X-1)*sqrt(x-1) =2
sqrt(x^2+3x-4)-x+1 =2
sqrt(x^2+3x-4)= x+1
square both sides
x^2+3x-4=(x+1)^2
x^2+3x-4=x^2+2x+1
x=4+1
x=5